# CamI2V: Camera-Controlled Image-to-Video Diffusion Model

### v1

@@ Abstract
Recently, camera pose, as a user-friendly and physics-related condition, has been introduced into text-to-video diffusion model for camera control. However, existing methods simply inject camera conditions through a side input. These approaches neglect the inherent physical knowledge of camera pose, resulting in imprecise camera control, inconsistencies, and also poor interpretability. In this paper, we emphasize the necessity of integrating explicit physical constraints into model design. Epipolar attention is proposed for modeling all cross-frame relationships from a novel perspective of noised condition. This ensures that features are aggregated from corresponding epipolar lines in all noised frames, overcoming the limitations of current attention mechanisms in tracking displaced features across frames, especially when features move significantly with the camera and become obscured by noise. Additionally, we introduce register tokens to handle cases without intersections between frames, commonly caused by rapid camera movements, dynamic objects, or occlusions. To support image-to-video, we propose the multiple guidance scale to allow for precise control for image, text, and camera, respectively. Furthermore, we establish a more robust and reproducible evaluation pipeline to solve the inaccuracy and instability of existing camera control measurement. We achieve a 25.5% improvement in camera controllability on RealEstate10K while maintaining strong generalization to out-of-domain images. Only 24GB and 12GB are required for training and inference, respectively. We plan to release checkpoints, along with training and evaluation codes. Dynamic videos are best viewed at https://zgctroy.github.io/CamI2V.
Figure 1: Rethinking condition in diffusion models. Diffusion models denoise along the gradient of log probability density function. At large noise levels, the high density region becomes the overlap of numerous noisy samples, resulting in visual blurriness. We point out that the effectiveness of a condition depends on how much uncertainty it reduces. From a new perspective, we categorize conditions into clean conditions (e.g. texts, camera extrinsics) that remain visible throughout the denoising process, and noised conditions (e.g. noised pixels in the current and other frames) whose deterministic information ${{\alpha}_{t}}x_{0}$ will be gradually dominated by the randomness of noise ${{\sigma}_{t}}\epsilon$.
@@ 1 Introduction
Condition or guidance (Ho & Salimans, 2022; Dhariwal & Nichol, 2021) is widely recognized as a crucial factor in enhancing generation quality. This is attributed to the fundamental principles that diffusion models denoise along the gradient of the log probability density function (score function) (Song et al., 2020), moving towards a high density region. However, this characteristic has varying effects at different noise levels (Tang et al., 2023a). As shown in Fig. 1(a), the high density region under high noise level becomes the overlap of numerous noisy samples, resulting in visual blurriness. By providing the model with conditions such as $c_{\text{dog}}$ and $c_{\text{cat}}$, it can rapidly eliminate incorrect generations. This illustrates that effective conditions can guide the model towards desired outcomes while reducing uncertainty.
@@
Consequently, incorporating physics-related or more detailed conditions into the diffusion model is an effective way of improving its world understanding. Considering that video generation requires providing condition for each frame, it is essential to identify a condition that is physics-related and also user-friendly. Recently, some camera-conditioned text-to-video diffusion models such as MotionCtrl (He et al., 2024a) and CameraCtrl (Wang et al., 2024d) have proposed using camera poses of each frame as a new type of condition. However, these methods simply inject camera conditions through a side input (like T2I-Adapter (Mou et al., 2024)) and neglect the inherent physical knowledge of camera pose, resulting in imprecise camera control, inconsistencies, and also poor interpretability. This raises a critical question: How can we effectively inject camera poses into the diffusion model to enhance its ability to physical world knowledge?
@@
In this paper, we integrate geometric priors into the model and propose a novel framework for injecting camera poses. We first identify shortcomings in existing attention mechanisms regarding the ability of tracking noised features across frames. As illustrated in Fig. 2, separated spatial and temporal attention serves as an indirect form of 3D attention. The cross-frame interaction in temporal attention is confined to features at the same location in the image, rendering it ineffective for tracking significant movements resulting from large camera shifts. 3D full attention is widely applied in advanced video diffusion models such as OpenSora (Zheng et al., 2024) and CogVideoX (Yang et al., 2024b), due to its extensive receptive field. From the novel perspective of the noised conditions mentioned in Fig. 1, the broad receptive field of 3D full attention allows it to access more noised conditions. However, noised conditions act as a double-edged sword; at high noise levels, deterministic information is often obscured by random noise, resulting in minimal reduction of uncertainty despite increased access. As previously highlighted in Fig. 1, the quality of a condition is determined solely by its ability to reduce the model’s uncertainty, rather than by the number of conditions it can access.
@@
To address these issues, we propose to apply plücker coordinates (Plücker, 1828) as absolute 3d ray embedding for implicit learning of 3d space and propose a novel epipolar attention mechanism that introduces an explicit constraint. This ensures that even when feature tracking is compromised under high noise conditions, only features along the epipolar line are aggregated. By doing so, our approach minimizes the search space and reduces potential errors, ultimately enhancing 3D consistency across frames and improving overall controllability. Our explicit epipolar attention can also be regarded as a form of clean condition, as it reliably filters out features that do not conform to 3D geometry across varying noise levels, significantly enhancing controllability and stability. Additionally, inspired by Timothée et al. (2024), we incorporate register tokens into epipolar attention to address scenarios where there are no intersections between frames, often caused by rapid camera movements, dynamic objects, or occlusions.
@@
  • We expand the camera-conditioned text-to-video diffusion model to support reference image at random frame and propose multiple guidance scale to allow for precise control for text, image, and camera, respectively.
  • From a novel perspective of noised condition, we propose epipolar attention added with register token as a explicit constraint for modeling all cross-frame relationships. This addresses the issue of noised features tracking due to the movement of camera, leading to precise camera control and better 3d consistency. It also reduces the controllability instability due to randomness introduced by noise during generation.
  • We point out and analyze the reasons for inaccurate measurement of camera controllability caused by the inherent limitations of SfM evaluator and re-establish a more robust, accurate and reproducible evaluation pipeline. Furthermore, we achieve the sota performance on RealEstate10K dataset and out-of-domain dataset.
@@ 2 Related Work
Camera-controlled Video Generation. AnimateDiff (Guo et al., 2023) utilizes LoRA (Hu et al., 2021) fine-tuning to achieve specific camera movements. MotionMaster (Hu et al., 2024) and Peekaboo (Jain et al., 2024) explore a training-free method for coarse-grained camera movement generation, but they lack precise control. VideoComposer (Wang et al., 2024b) offers global motion guidance by adjusting pixel-level motion vectors. In contrast, MotionCtrl (Wang et al., 2024d), CameraCtrl (He et al., 2024a), and Direct-a-Video (Yang et al., 2024a) incorporate camera pose information as side input; however, these methods primarily focus on text-to-video generation and do not effectively leverage 3D geometric priors in camera pose. Concurrent work, CamCo (Xu et al., 2024a), also facilitates controllable camera generation in the image-to-video task by using epipolar lines to ensure consistency between generated frames and a single reference frame only. However, it does not account for scenarios where frames lack overlapping regions with the reference frame and can thus be regarded as a degenerate version of our approach.
@@ 3.1 Preliminaries
3D Ray Embedding. Considering camera intrinsics $K\in{\mathbb{R}}^{3\times 3}$ and extrinsics (rotation $R\in\text{SO}(3)$, translation $T\in{\mathbb{R}}^{3}$), it parameterizes the transform from world coordinates to pixel coordinates by projection $u=K\begin{bmatrix}R~{}|~{}T\end{bmatrix}x$. This low-dimensional representation may hinder neural networks from direct regression. Instead, we follow (Tseng et al., 2023) to represent cameras as ray bundles:
@@
$\mathcal{L}=\mathbb{E}_{z,c_{\text{txt}},c_{\text{img}},\epsilon\sim\mathcal{N}(0,\mathbf{\mathrm{I}}),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(z_{t},c_{\text{txt}},c_{\text{img}},t\right)\right\|_{2}^{2}\right],$ (4)
where $z\in\mathbb{R}^{F\times H\times W\times C}$ is the latent code of video data with $F,H,W,C$ being frame, height, width, and channel. Besides, $c_{\text{text}}$ is the text prompt for input video, and $c_{\text{img}}$ is the reference frame of video. A noise-corrupted latent code $z_{t}$ from the ground-truth $z_{0}$ is formulated as $z_{t}=\alpha_{t}z_{0}+\sigma_{t}\epsilon$, where $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$, $\alpha_{t}$ and $\sigma_{t}$ are hyperparameters to control the diffusion process.
Figure 4: Pipeline of camera-conditioned image-to-video diffusion model.
@@ 3.3 Epipolar Attention for Noised Features Tracking
where $F_{ij}$ is the camera fundamental matrix of two frames, which can be derived as $F_{ij}=K_{j}^{\rm-T}\cdot E_{ij}\cdot K_{i}^{-1}$ given the camera intrinsics $K_{i},K_{j}\in{\mathbb{R}}^{3\times 3}$ and the camera essential matrix $E_{ij}\in{\mathbb{R}}^{3\times 3}$. We transform the camera pose of the $j$-th frame to be relative to the $i$-th frame for simplicity, thus it holds that $E_{ij}=T_{i\rightarrow j}\times R_{i\rightarrow j}$, where $R_{i\rightarrow j}\in{\mathbb{R}}^{3\times 3}$ and $T_{i\rightarrow j}\in{\mathbb{R}}^{3}$ are the relative rotation matrix and translation vector, respectively. Due to the contiguous representation of the epipolar line $l_{ij}=Ax+By+C$, we convert it to attention mask by calculating per-pixel distance at coordinate $(u^{\prime},v^{\prime})$ on the $j$-th frame to the epipolar line as
@@
Epipolar attention. We extend current temporal attention with epipolar constraint to leverage cross-frame relationship and inject geometry consistency for video generation. We denote the noised latent of the $i$-th frame at timestamp $t$ as $z_{t}^{i}\in\mathbb{R}^{hw\times c}$. And the query, key and value are defined as $q=W_{q}z_{t}^{i}\in\mathbb{R}^{hw\times c}$, $k=W_{k}\{z_{t}^{j}\}_{j=1}^{N}\in\mathbb{R}^{Nhw\times c}$ and $v=W_{v}\{z_{t}^{j}\}_{j=1}^{N}\in\mathbb{R}^{Nhw\times c}$, where $W_{q}$, $W_{k}$ and $W_{v}$ are the projection matrices respectively. Given the epipolar attention mask $m_{i}\in\mathbb{R}^{hw\times Nhw}$ introduced in Section 3.3, our epipolar attention that captures relevant contextual information between the $i$-th frame and all $N$ frames is then computed as
${\rm EpipolarAttn}(z_{t}^{i},\{z_{t}^{j}\}_{j=1}^{N},m_{i})=\mathop{\rm softmax}\left(\frac{qk^{\rm T}}{\sqrt{d}}\odot m_{i}\right)v,$ (7)
@@
Register tokens for special cases. For videos with significant camera movements, dynamic objects, or occlusions, there may be cases where some pixels from the $i$-th frame have no corresponding epipolar lines within the image planes of all $N$ frames. This situation can lead to a zero epipolar mask, affecting the computational stability of the epipolar attention mechanism.
To address this issue, we draw inspiration from Timothée et al. (2024) and introduce additional register tokens to the input sequence as a straightforward solution, as illustrated in Fig. 6. Additionally, register tokens are learnable, enabling adaptive learning to address various special cases.
@@ 4 Metrics and Evaluation
Before calculating metrics, we canonicalize the estimated camera-to-world matrices by converting each frame relative to the first frame and normalizing the scene scale using the ${\cal L}_{2}$ distance from the first camera to the furthest cameras. To account for randomness introduced by GLOMAP, we conduct five individual trials for each of the 1,000 sampled videos, averaging only those trials that are successful per sample. The final metrics, including $\rm RotErr$, $\rm TransErr$, and $\rm CamMC$, are averaged on a sample-wise basis.
RotError. We evaluate per-frame camera-to-world rotation accuracy by the relative angles between ground truth rotations $R_{i}$ and estimated rotations $\tilde{R}_{i}$ of generated frames. We report accumulated rotation error along 16 frames in radians.
@@
TransError. We evaluate per-frame camera trajectory accuracy by the camera location in the world coordinate system, i.e. the translation component of camera-to-world matrices. We report the sum of ${\cal L}_{2}$ distance between ground truth translations $T_{i}$ and generated translations $\tilde{T}_{i}$ for all 16 frames.
@@
CamMC. We also evaluate camera pose accuracy by directly calculating ${\cal L}_{2}$ similarity of per-frame rotations and translations as a whole. We sum up the results of 16 frames.
@@
FVD. Additionally, to ensure that proposed method coherently improve generative capability and visual quality of base I2V model, we evaluate the distance of generated frames from training distribution by Fréchet Video Distance (FVD) (Unterthiner et al., 2018) [^1] We use “fvdcal” to calculate FVD metric..

### v3

@@ Abstract
Recent advancements have integrated camera pose as a user-friendly and physics-informed condition in video diffusion models, enabling precise camera control. In this paper, we identify one of the key challenges as effectively modeling noisy cross-frame interactions to enhance geometry consistency and camera controllability. We innovatively associate the quality of a condition with its ability to reduce uncertainty and interpret noisy cross-frame features as a form of noisy condition. Recognizing that noisy conditions provide deterministic information while also introducing randomness and potential misguidance due to added noise, we propose applying epipolar attention to only aggregate features along corresponding epipolar lines, thereby accessing an optimal amount of noisy conditions. Additionally, we address scenarios where epipolar lines disappear, commonly caused by rapid camera movements, dynamic objects, or occlusions, ensuring robust performance in diverse environments. Furthermore, we develop a more robust and reproducible evaluation pipeline to address the inaccuracies and instabilities of existing camera control metrics. Our method achieves a 25.64% improvement in camera controllability on the RealEstate10K dataset without compromising dynamics or generation quality and demonstrates strong generalization to out-of-domain images. Training and inference require only 24GB and 12GB of memory, respectively, for 16-frame sequences at 256×256 resolution. We will release all checkpoints, along with training and evaluation code. Dynamic videos are best viewed at https://zgctroy.github.io/CamI2V
Figure 1: Rethinking condition in diffusion models. Diffusion models denoise along the gradient of log probability density function. At large noise levels, the high density region becomes the overlap of numerous noisy samples, resulting in visual blurriness. We point out that the effectiveness of a condition depends on how much uncertainty it reduces. From a new perspective, we categorize conditions into clean conditions (e.g. texts, camera extrinsics) that remain visible throughout the denoising process, and noisy conditions (e.g. noised pixels in the current and other frames) whose deterministic information ${{\alpha}_{t}}x_{0}$ will be gradually dominated by the randomness of noise ${{\sigma}_{t}}\epsilon$.
@@ 1 Introduction
Condition or guidance (Ho & Salimans, 2022; Dhariwal & Nichol, 2021) is widely recognized as a crucial factor in enhancing generation quality. This is attributed to the fundamental principles that diffusion models denoise along the gradient of the log probability density function (score function) (Song et al., 2020), moving towards a high density region. However, this characteristic has varying effects at different noise levels (Tang et al., 2023a). As shown in Fig. 1(a), the high density region under high noise level becomes the overlap of numerous noisy samples, resulting in visual blurriness. By providing the model with conditions such as $c_{\text{dog}}$ and $c_{\text{cat}}$, it can rapidly eliminate incorrect generations. This illustrates that adding more conditions can guide the model towards desired outcomes while reducing uncertainty.
@@
Consequently, incorporating physics-related or more detailed conditions into the diffusion model is an effective way of improving its world understanding. Considering that video generation requires providing condition for each frame, it is essential to identify a condition that is physics-related but also user-friendly. Recently, some camera-conditioned text-to-video diffusion models such as MotionCtrl (He et al., 2024a) and CameraCtrl (Wang et al., 2024d) have proposed using camera poses of each frame as a new type of condition. However, these methods simply inject camera conditions through a side input (like T2I-Adapter (Mou et al., 2024)) and neglect the inherent physical knowledge of camera pose, resulting in imprecise camera control, inconsistencies, and also poor interpretability.
@@
In this paper, we identify one of the key challenges of camera-controlled image-to-video diffusion models as how to effectively model noisy cross-frame interactions to enhance geometry consistency and camera controllability. As illustrated in Fig. 2, separated spatial and temporal attention serves as an indirect form of 3D attention. The cross-frame interaction in temporal attention is confined to features at the same location in the image, rendering it ineffective for tracking significant movements resulting from large camera shifts. 3D full attention is widely applied in advanced video diffusion models such as OpenSora (Zheng et al., 2024) and CogVideoX (Yang et al., 2024b), due to its extensive receptive field. From the novel perspective of the noisy conditions mentioned in Fig. 1, the broad receptive field of 3D full attention allows it to access more noisy conditions. However, we argue that accessing more noisy conditions does not necessarily reduce uncertainty and thus not necessarily lead to better performance due to the randomness inherent in the noise. As previously highlighted in Fig. 1, the quality of a condition is determined by its ability to reduce the model’s uncertainty, rather than its quantity.
@@
To address these issues, we have found that applying epipolar constraints is one of the most suitable way to prevent the model from being misled by noise. By restricting attention to features along the epipolar lines, the model can interact with more relevant and less noisy information, improving cross-frame interactions in diffusion models. Specifically, we propose to apply Plücker coordinates (Plücker, 1828) as absolute 3D ray embedding for implicit learning of 3D space and propose a epipolar attention mechanism that introduces an explicit constraint. By doing so, our approach minimizes the search space and reduces potential errors, ultimately enhancing 3D consistency across frames and improving overall controllability. Additionally, inspired by Timothée et al. (2024), we incorporate register tokens into epipolar attention to address scenarios where there are no intersections between frames, often caused by rapid camera movements, dynamic objects, or occlusions.
@@
  • We identify one of the key challenges of camera-controlled image-to-video diffusion models as effectively modeling noisy cross-frame interactions to enhance geometry consistency and camera controllability.
  • Well-motivated by the relationship between the quality of a condition and its ability to reduce uncertainty, we innovatively interpret noisy cross-frame features as a form of noisy condition and propose to apply epipolar attention to access an optimal amount of noisy condition. We also address scenarios where epipolar lines disappear by register tokens.
  • We point out and analyze the reasons for inaccurate measurement of camera controllability caused by the inherent limitations of SfM evaluator and re-establish a more robust, accurate and reproducible evaluation pipeline. We achieve a 32.96%, 25.64%, 20.77% improvement over CameraCtrl on RotErr, CamMC, TransErr on the RealEstate10K dataset without compromising dynamics, generation quality, or generalization on out-of-domain images.
@@ 2 Related Work
Camera-controlled Video Generation. AnimateDiff (Guo et al., 2023) utilizes LoRA (Hu et al., 2021) fine-tuning to achieve specific camera movements. MotionMaster (Hu et al., 2024) and Peekaboo (Jain et al., 2024) explore a training-free method for coarse-grained camera movement generation, but they lack precise control. VideoComposer (Wang et al., 2024b) offers global motion guidance by adjusting pixel-level motion vectors. In contrast, MotionCtrl (Wang et al., 2024d), CameraCtrl (He et al., 2024a), and Direct-a-Video (Yang et al., 2024a) incorporate camera pose information as side input; however, these methods primarily focus on text-to-video generation and do not effectively leverage 3D geometric priors in camera pose. CamCo (Xu et al., 2024a) also facilitates controllable camera generation in the image-to-video task by using epipolar attention (Kant et al., 2024; Tseng et al., 2023) to ensure consistency between generated frames and a single reference frame only. However, it does not account for scenarios where frames lack overlapping regions with the reference frame and can thus be regarded as a degenerate version of our approach.
@@ 3.1 Preliminaries
3D Ray Embedding. We follow CameraCtrl (He et al., 2024a) to apply plücker embedding as global positional embedding. Considering camera intrinsics $K\in{\mathbb{R}}^{3\times 3}$ and extrinsics (rotation $R\in\text{SO}(3)$, translation $T\in{\mathbb{R}}^{3}$), it parameterizes the transform from world coordinates to pixel coordinates by projection $u=K\begin{bmatrix}R~{}|~{}T\end{bmatrix}x$. This low-dimensional representation may hinder neural networks from direct regression. Instead, we follow  (Tseng et al., 2023) to represent cameras as ray bundles:
@@
$\mathcal{L}=\mathbb{E}_{z,c_{\text{txt}},c_{\text{img}},\epsilon\sim\mathcal{N}(0,\mathbf{\mathrm{I}}),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},c_{\text{txt}},c_{\text{img}},t\right)\right\|_{2}^{2}\right],$ (4)
where $\mathbf{z}\in\mathbb{R}^{F\times H\times W\times C}$ is the latent code of video data with $F,H,W,C$ being frame, height, width, and channel. Besides, $c_{\text{text}}$ is the text prompt for input video, and $c_{\text{img}}$ is the reference frame of video. A noise-corrupted latent code $\mathbf{z}_{t}$ from the ground-truth $z_{0}$ is formulated as $\mathbf{z}_{t}=\alpha_{t}z_{0}+\sigma_{t}\epsilon$, where $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$, $\alpha_{t}$ and $\sigma_{t}$ are hyperparameters to control the diffusion process.
Figure 4: Pipeline of camera-controlled image-to-video diffusion model. We follow CameraCtrl to add a learnable pose encoder and a linear projection to process plucker embeddings as a global positional embedding. Epipolar attention is added between spatial and temporal attention.
@@ 3.3 Epipolar Attention for Noised Features Tracking
where $F_{ij}$ is the camera fundamental matrix of two frames, which can be derived as $F_{ij}=K_{j}^{\rm-T}\cdot E_{ij}\cdot K_{i}^{-1}$ given the camera intrinsics $K_{i},K_{j}\in{\mathbb{R}}^{3\times 3}$ and the camera essential matrix $E_{ij}\in{\mathbb{R}}^{3\times 3}$. We transform the camera pose of the $j$-th frame to be relative to the $i$-th frame for simplicity, thus it holds that $E_{ij}=T_{i\rightarrow j}\times R_{i\rightarrow j}$, where $R_{i\rightarrow j}\in{\mathbb{R}}^{3\times 3}$ and $T_{i\rightarrow j}\in{\mathbb{R}}^{3}$ are the relative rotation matrix and translation vector, respectively. Due to the contiguous representation of the epipolar line $l_{ij}=Ax+By+C$, we convert it to attention mask by calculating per-pixel distance $D$ at coordinate $(u^{\prime},v^{\prime})$ on the $j$-th frame to the epipolar line as
@@
Epipolar attention. We extend current temporal attention with epipolar constraint to leverage cross-frame relationship and inject geometry consistency for video generation.
We denote the query, key and value as $q\in\mathbb{R}^{hw\times c}$, $k\in\mathbb{R}^{Nhw\times c}$ and $v\in\mathbb{R}^{Nhw\times c}$, respectively. Given the epipolar attention mask $m\in\mathbb{R}^{hw\times Nhw}$ introduced in Section 3.3, our epipolar attention that captures relevant contextual information between the $i$-th frame and all $N$ frames is then computed as
${\rm EpipolarAttn}(q,k,v,m)=\mathop{\rm softmax}\left(\frac{qk^{\rm T}}{\sqrt{d}}\odot m\right)v,$ (7)
@@
Register tokens for scenarios where epipolar lines disappear. For videos with significant camera movements, dynamic objects, or occlusions, there may be cases where some pixels from the $i$-th frame have no corresponding epipolar lines within the image planes of all $N$ frames. This situation can lead to a zero epipolar mask, affecting the computational stability of the epipolar attention mechanism.
To address this issue, we draw inspiration from Timothée et al. (2024) and introduce additional register tokens to the input sequence as a straightforward solution, as illustrated in Fig. 6. Additionally, register tokens are learnable, enabling adaptive learning to address various special cases. Without register tokens to serve as placeholders, we may encounter the zero length of key/value tokens and fail to calculate attention
@@ 4 Metrics and Evaluation
Before calculating metrics, we canonicalize the estimated camera-to-world matrices by converting each frame relative to the first frame and normalizing the scene scale using the ${\cal L}_{2}$ distance from the first camera to the furthest cameras. To account for randomness introduced by GLOMAP, we conduct five individual trials for each of the 1,000 sampled videos, averaging only those trials that are successful per sample. The final metrics, including RotError, TransError, and CamMC, are averaged on a sample-wise basis.
RotError (He et al., 2024a). We evaluate per-frame camera-to-world rotation accuracy by the relative angles between ground truth rotations $R_{i}$ and estimated rotations $\tilde{R}_{i}$ of generated frames. We report accumulated rotation error along 16 frames in radians.
@@
TransError (He et al., 2024a). We evaluate per-frame camera trajectory accuracy by the camera location in the world coordinate system, i.e. the translation component of camera-to-world matrices. We report the sum of ${\cal L}_{2}$ distance between ground truth translations $T_{i}$ and generated translations $\tilde{T}_{i}$ for all 16 frames.
@@
CamMC (Wang et al., 2024d). We also evaluate camera pose accuracy by directly calculating ${\cal L}_{2}$ similarity of per-frame rotations and translations as a whole. We sum up the results of 16 frames.
@@
FVD (Unterthiner et al., 2018). Additionally, to ensure that proposed method coherently improve generative capability and visual quality of base I2V model, we evaluate the distance of generated frames from training distribution by Fréchet Video Distance (FVD).